This is the Measurement of Chloride

Water Sample Initial volume of water (mL) Average volume of AgNO3 added (mL)
Lake water 100.0 14.62
Wastewater effluent 25.0 20.31
Tap water 100.0 13.03

Materials

-Water samples (Lake water, Wastewater effluent, and Tap water)

-0.010M AgNO3

-0.25 M Na2CrO4

All of the calculation need to be completed for each of the three water samples.

1. Moles of AgNO3 used ot react Cl (Hint: Use the concentration of the AgNo3 solution (in M) and the average volume added (in L)

2. Moles of Cl reacted (Hint: the chemical equation for the reaction of chloride with silver nitrate shows a 1:1 molar ratio between AgNO3 and Cl: AgNO3 +Cl- -> AgCl +NO3-)

3. Concentration of Cl (in units M, moles/L) (Hint: Use your calculated moles of Cl and the volume of the water sample (in L) used in the titration)
Answer #1

1.moles of AgNO3=volume*molarity

molarity of AgNO3=0.01M

FOR LAKE WATER:volume of AgNO3=14.62mL

Then mole=14.62*10-3*0.01=0.0001462

FOR WASTE WATER EFFLUENT:volume=20.31ml=20.31*10-3 L

Then mole=20.31*10-3 L*0.01M=0.0002031

FOR TAP WATER:volume=13.03*10-3L

mole=13.03*10-3L*0.01=0.0001303

2.In that above equation 1mol of AgNO3 reacts with 1 mole of chloride ion.So in each case mole of chloride reacted is same as mole of AgNO3 is used .

FOR LAKE WATER:mole of Cl- reacted=0.0001462

FOR WASTE WATER EFFLUENT:mole of Cl- reacted=0.0002031

FOR TAP WATER:mole of Cl- reacted=0.0001303

3. concentration of Cl-=mole/volume of sample water taken initially

FOR LAKE WATER:volume of sample water=100mL=0.1L

concentration of Cl-=0.0001462/0.1=0.001462M

FOR WASTE WATER EFFLUENT:volume of sample water=25mL=0.025L

concentration of Cl-=0.0002031/0.025L=0.008124M

FOR TAP WATER:volume of sample water=100mL=0.1L

concentration of Cl-=0.0001303/0.1L=0.001303M